At the state fair, Felipe bought 2 tshirts and two soft drinks for $44.00. Heather bought one tshirt and three soft drinks and paid $30. How much does a tshirt cost? How much is a soft drink?
1
Expert's answer
2012-11-14T07:25:58-0500
2x+2y=44
X+3y=30
X=30-3y
60-6y+2y=44 & -4y=-16 y=4
X=30-3*4=30-12=18
x-t-shirts
y-soft drinks
Answer: t-shirts 18$, soft drinks 4$